OpenStudy (anonymous):

If you had a 5 digit pin (numbers 0 - 9) with all combinations of digits being equally likely, what is the probability of guessing the pin if you try 3 times with a different pin each time?

OpenStudy (anonymous):

I was thinking you would do 1 1 1 ----------- + ----------- + ---------- 10^5 10^5 - 1 10^5 - 2 But I'm not completely sure

OpenStudy (anonymous):

since there are 10^5 total combinations and you use up one each time

OpenStudy (anonymous):

\[\frac{ 10! }{5!*5!} = 36\]

OpenStudy (anonymous):

There are 36 Combination

OpenStudy (anonymous):

don't you have 10 different numbers to choose for each digit so 10*10*10*10*10 combinations?

OpenStudy (anonymous):

\[\binom{10}{5} = \frac{10!}{5!*5!}\]

hartnn (hartnn):

what?!! @Yahoo! u have 3 trial, in first case, if u guess it correctly, its 1/(10^5) in 2nd trial, U HAVE NOT GUESSED IT CORRECTLY, so (1-1/10^5)*(1/(10^5-1))

OpenStudy (anonymous):

U have 10 Digit and u r taking 5 at a time

hartnn (hartnn):

in third trial , u have not guessed it correctly both times , so (1-1/10^5)*(1-1/(10^5-1)*(1/(10^5-2)))

hartnn (hartnn):

add all these 3

OpenStudy (anonymous):

where does this come from (1-1/10^5)?

hartnn (hartnn):

u will try for 2nd time only if your first trial was unsuccessful so unsuccessful = 1-1/10^5

hartnn (hartnn):

to guess it correct 2nd time u need to multiply that by 1/(10^5-1)

OpenStudy (anonymous):

what is that 1 from in the 1-(1/10^5)?

hartnn (hartnn):

(1st guess correct) OR (1st guess incorrect AND 2nd guess correct) OR (1st guess incorrect AND 2nd guess incorrect AND 3rd guess correct)

hartnn (hartnn):

correct = p, then not correct = 1-p

OpenStudy (anonymous):

what is the p?

hartnn (hartnn):

but r u sure about 10^5

hartnn (hartnn):

probability of guessing correct

OpenStudy (anonymous):

I am pretty sure. There was a similar example in my text

hartnn (hartnn):

p=1/10^5

OpenStudy (anonymous):

what if you just assume the second one is your first guess with just 1 less answer and just find the probability of 1/(10^5 - 1)

hartnn (hartnn):

but why would u take a 2nd guess if your 1st one is correct?

OpenStudy (anonymous):

the question doesn't say that you stop when you get the right pin, only that you make 3 guesses

OpenStudy (anonymous):

they just want to know, of the 3 guesses you made, how likely is it that one of those choices would be the correct pin

hartnn (hartnn):

3 cases (correct guess in 1st try itself) OR (correct guess in 2nd try only) OR (correct guess in 3rd try only) which comes out to be (1st guess correct) OR (1st guess incorrect AND 2nd guess correct) OR (1st guess incorrect AND 2nd guess incorrect AND 3rd guess correct)

hartnn (hartnn):

notice 'one of the..'

hartnn (hartnn):

only one must be correct

hartnn (hartnn):

so for 2nd guess, 1st must be incorrect

OpenStudy (anonymous):

well only 1 must be able to be correct since you need to choose different digits each time

OpenStudy (anonymous):

assume that you make randomly make 3 different pins. What is the probability of 1 of those 3 different pins being the correct one?

hartnn (hartnn):

1/3?

OpenStudy (anonymous):

well there is no guarantee any of them are correct

OpenStudy (anonymous):

3 choices out of 10^5 possible combinations

OpenStudy (anonymous):

and each time you have one less option because they all must be different

hartnn (hartnn):

i would still stick to my (1-...) thingy different, thats why 10^5 in first denom, 10^5-1 in 2nd denom and 10^5-2 in 3rd denom...

OpenStudy (anonymous):

I just don't really understand that 1- thingy

OpenStudy (anonymous):

if it's there because you want to stop and check after each guess I don't think the question is asking to do that

hartnn (hartnn):

(1st guess correct) OR (1st guess incorrect AND 2nd guess correct) OR (1st guess incorrect AND 2nd guess incorrect AND 3rd guess correct) u didn't understand this, right ?

OpenStudy (anonymous):

yeah ok I understand that part. Just not what you are putting it into cases at all. As far as I can see it want you to see what the probability is that of 3 randomly chosen pins, one of them matches up to the correct pin

hartnn (hartnn):

ok,ok....when i solved such problems......i always did this....if the question is really just about guessing , then u are correct.....sorry for wasting your time...

OpenStudy (anonymous):

no no, I appreciate your input, it really helps me find the correct solution.

OpenStudy (anonymous):

I was never taught the 1- thingie yet so I don't think they would put it on an assignment as well